In the pair of supply and demand equations below, x represents the quantity demanded in units of a thousand and p the unit price in dollars, find the equilibrium quantity and the equilibrium price.
p = 108 − 4x^2 and p = x^2 + 10x + 33

equilibrium quantity ??? Thousand units

equilibrium price ??? Dollars

Answer by stanbon(75887)

In the pair of supply and demand equations below, x represents the quantity demanded in units of a thousand and p the unit price in dollars, find the equilibrium quantity and the equilibrium price.
p = 108 − 4x^2 and p = x^2 + 10x + 33
-------
Use substitution to solve::
108-4x^2 = x^2 + 10x + 33
5x^2 + 10x -75 = 0
x^2 + 2x - 15 = 0
-----
(x-5)(x+3) = 0
Positive solution::
x = 5
--------
equilibrium quantity ??? 5 Thousand units
equilibrium price ??? 108-4*5^2 = 108-100 = $8 Dollars
-----------------
Cheers,
Stan H.
-------------